OpenStudy (anonymous):

The sum of the angle measures of any triangle is 180. the second angle measures three times first, and the 3rd angle measures 20 less than the first. what is the measure of each of the angle?

OpenStudy (anonymous):

We have a set of 3 equations, given angle 1, a1, angle 2, a2, angle 3, a3 a1 = a1 a2 = 3a1 a3 = a1/20 a1 + a2 + a3 = 180 Set the last equation to only have one variable using substitution a1 + 3a1 + a1/20 = 180 79a1/ 20 = 180 79 a1 = 3600 a1 = 45.6 degress Now we know a1 and can solve for all the other angles easile a2 = 3a1 a2 = 3(45.6) a2 = 136.8 a3 = a1/20 a3 = 45.6/2- a3 = 4.56/2 a3 = 2.28

OpenStudy (anonymous):

how did u get 79?

OpenStudy (anonymous):

nvm thanks